A IS three time old as b .C was twice as old as A four year ago .in four year time A will 31. What the present age B and C?
Answer by ankor@dixie-net.com(22740)

Write an equation for each statement
:
A IS three time old as B
A = 3B
;
C was twice as old as A four year ago
C -4 = 2(A-4)
:
in four year time A will 31.
A + 4 = 31
A =31- 4
A = 27 yrs, his present age
:
Then using the 2nd equation
c - 4 = 2(27 - 4)
C - 4 = 2(23)
c = 46 + 4
C = 50 yrs, his present age
:
and the 1st equation
27 = 3B
Divide both sides by 3
9 = B
:
What the present age B and C?
B is 9, C is 50